A remotely located air sampling station can be powered by solar cells or by running an electric line to the site and using conventional power. Solar cells will cost $11,000 to install and will have a useful life of 4 years with no salvage value. Annual costs for inspection, cleaning, etc. are expected to be $2,000. A new power line will cost $18,000 to install, with power costs expected to be $1,000 per year. Since the air sampling project will end in 4 years, the salvage value of the line is considered to be zero. At an interest rate of 8% per year, which alternative should be selected on the basis of a future worth analysis? The future worth of solar cells is $ Not attempted and that of electric line is $ Not attempted .
Answer #1
FV of the solar cells = 11000*1.08^4+2000*(1.08^4-1)/0.08 = $        23,977.60
FV of the electric line = 18000*1.08^4+1000*(1.08^4-1)/0.08 = $        28,994.91
The solar alternative should be selected as it has lower FW Cost.